# Triple Phase Transitions: Understanding the Learning Dynamics of Large Language Models from a Neuroscience Perspective

### v1

@@ 1 Introduction
Figure 1: Overview of the study. a Brain encoding analysis. b Probing analysis (top) and the benchmark analysis (bottom). c Three phase-transition phenomena during the learning process of LLMs, as identified through the results of the encoding, probing, and benchmark analyses. Red, green, and blue lines indicate encoding, probing, and benchmark accuracies, respectively, accros all LLM checkpoints.
@@ 3.1 Large Language Models
We analyze the learning dynamics of LLMs from three distinct perspectives. For this, we use four pre-trained models with available training checkpoints: OLMo-2 (OLMo et al., 2024), OLMo-0724 (Groeneveld et al., 2024), and LLM-jp (LLM-jp et al., 2024) for the main analysis, and Amber (Liu et al., 2023) for the control analysis. Table 1 presents an overview of these LLMs. OLMo-2, OLMo-0724, and Amber are English-centric LLMs trained on publicly accessible datasets, while LLM-jp is a bilingual Japanese-English model trained on a roughly equal mix of Japanese and English. Each model uses a different training corpus and tokenizer, with vocabulary sizes of 100,352 (OLMo-2), 50,304 (OLMo-0724), 99,584 (LLM-jp), and 32,000 (Amber). These LLMs have between 6.74–7.3B parameters, 32 hidden layers, and a hidden dimension of 4,096. They are all based on a decoder-only Transformer architecture (Vaswani et al., 2017), but each LLM incorporates a few critical modifications. There are also notable differences across these LLMs in terms of their layer normalization and attention mechanisms (see Section A.1 for architectural details). The multilayer perceptron (MLP) layers require most parameters and represent essential features (Bereska & Gavves, 2024; Geva et al., 2021). Thus, we use their neural activation as the internal representations of each model. We use 28 checkpoints for OLMo-2 (1B-3,896B training tokens), 23 for OLMo-0724 (4B–2,724B), 27 for LLM-jp (4.2M–1,258B), and 18 for Amber (3.5B–1,259B). See Section A.1 for further details on the model sources and specific checkpoints.
@@ 3.2 Brain Encoding Models
Our initial approach for analyzing the learning dynamics of the LLMs involves an investigation of how their activations progressively align with brain activity during the training process. Specifically, for each checkpoint described in Section 3.1, we perform an encoding analysis by evaluating the prediction accuracy of a learned linear mapping from each layer’s activations to brain activity (Naselaris et al., 2011; Nishimoto et al., 2011; Huth et al., 2012) (Figure 1a). These analysis are conducted separately for each participant.
@@
We use the Narrative Movie fMRI Dataset (Yamaguchi et al., 2024; Nakagi et al., 2024) for our brain encoding analysis. This dataset provides brain activity data from six healthy participants with normal or corrected-normal vision (three females; ages 22–40, mean = 28.7), while they freely watched 8.3 hours of movies or drama series inside a 3T functional Magnetic Resonance Imaging (fMRI) scanner. All participants were right-handed and native Japanese speakers. The dataset includes nine video clips of movies or drama series as stimuli: eight international titles and one Japanese title. The international titles were dubbed into Japanese, allowing the participants to view all clips in Japanese. The dataset also contains three types of natural language annotation from the videos. We use the Narrative Content annotation, which describes the background story of the scene at 5-s intervals, for the main analysis and the Objective Information annotation, which describes the objects in the scene every second, for the control analysis, both in English and Japanese. Section A.2 provides additional details regarding the dataset. In this study, we use 29,993 seconds of data covering all six participants. We divide the data into training and test datasets. All prediction performance results are computed using the test dataset. Specifically, we use the fMRI scanning sessions corresponding to the last split of each movie or drama series (7,737 seconds in total) as the test data. The remainder of the sessions (22,262 seconds in total) constitutes the training data.
@@
We extract the activations of the LLMs for the annotations from each hidden layer. They consist of several tokens for each time point. We then average the activations across tokens. Because multiple annotations exist for each second, we extract the activations for each annotator and then average them across all annotators. We then train an L2-regularized linear model to predict the voxel-level brain activity from the corresponding activations of the LLM neurons. We estimate the model weights from the training data, then apply them to the test data. The regularization parameters are determined for each voxel by cross-validation during training. We then evaluate the prediction accuracy by computing the Pearson’s correlation coefficients between the predicted and measured fMRI signals. Statistical significance is assessed using a blockwise permutation test that compares the correlations between predicted and measured signals against the correlations obtained after shuffling the measured signals. We set the threshold for statistical significance to $p<0.05$ and correct for multiple comparisons using the FDR procedure. We model the hemodynamic delay in the BOLD signal, assumed to be 8–10 seconds. See Section A.3 for additional details about model construction and region-of-interest (ROI) analyses. In the main analysis, we focus on later layers within each LLM. Specifically, we focus on layer 25 in both OLMo-2 and LLM-jp, and layer 30 in OLMo-0724. These layers show the most substantial checkpoint-wise changes and are strongly involved in the emergent behavior of each model, as verified in all three evaluations (encoding, probing, and benchmark analyses; see Section A.6 for details). We confirm that the layers surrounding the main focal layer exhibit similar patterns of results (see Figure B.5).
@@ 3.3 Probing
As the second approach, we investigate how downstream task-relevant representations are progressively acquired within the LLMs during the training process. To this end, we perform a probing analysis by evaluating how well linear models can predict neuron activations from the answer labels of downstream tasks across all hidden layers (Figure 1b, top). These analysis are conducted separately for each LLM checkpoint.
@@
We use four downstream tasks: Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2020), CommonsenseQA (CSQA) (Talmor et al., 2019), AI2 Reasoning Challenge (ARC) (Clark et al., 2018), and HellaSwag (Zellers et al., 2019). MMLU assesses broad knowledge and problem-solving abilities using multidisciplinary coverage of 57 subjects, CSQA tests everyday conceptual commonsense reasoning, ARC probes elementary-level scientific knowledge, and HellaSwag assesses contextual commonsense reasoning in typical scenarios. For both the probing and benchmark analyses (see Section 3.4), we use 5-shot prompts in both English and Japanese. We exclude prompts exceeding each model’s maximum context length. See Section A.4 for additional details about downstream datasets. In the probing analysis, each dataset is split into training and test datasets at a 4:1 ratio. Because MMLU comprises multiple subject areas, we split the dataset by subject. All prediction performance results are computed using the test dataset. Examples of the prompts used in these datasets are presented in Section A.7.
@@
We first feed the downstream task prompts into the LLMs and extract the final-token neuron activations in each layer. For each sample in the downstream tasks, we create an answer matrix (the number of samples $\times$ the number of choices), where the correct and incorrect choices are labeled as 1 and 0, respectively. Our objective is to learn a mapping from this answer matrix to the neuron activations across all layers of the LLMs. To this end, we use L2-regularized linear regression to estimate the linear weights that transform the answer matrix corresponding to the training data into the observed neuron activations in the LLMs, and then apply these learned weights to the test data. The regularization parameter is optimized via 4-fold cross-validation for each neuron in the training dataset. To mitigate the subject-based bias of MMLU, we shuffle the training indices, and then perform cross-validation to ensure balanced distributions in each fold. We then evaluate the prediction accuracy by computing the Pearson’s correlation coefficients between the predicted and actual activations. The amount of information necessary for each task that is retained at each LLM checkpoint is quantitatively evaluated at the neuron level. In this study, we focus on the same layers as in the encoding analysis (Section 3.2).
@@ 4.1 Learning Dynamics of LLMs
Figure 3: Changes in the relationship with the brain. Differences in encoding accuracy among checkpoints for three participants (DM06, DM03, and DM07), projected onto the inflated (top, lateral, and medial views) and flattened cortical surface (occipital areas are at the center, only for DM06), for both the left and right hemispheres. We have chosen OLMo-2’s checkpoints that capture the transitions at each phase of the learning dynamics. Brain regions with significant accuracy at either of the two checkpoints are colored ($p<0.05$, FDR-corrected). Voxels that exhibit higher accuracy at the later checkpoint are indicated in red, whereas those that exhibit higher accuracy at the earlier checkpoint are indicated in blue.
@@ 4.2 What Happens During Each Phase?
Overall, these findings indicate that once the LLMs begin following the task instructions (Phase 1), their internal representations exhibit better global alignment with brain activity. During the subsequent stagnation phase (Phase 2), this global alignment partially recedes, and after Phase 3, the LLMs realign with the brain, reinforcing the similarity to brain regions involved in semantic processing, in particular. Notably, these patterns only emerge when the input language matches one on which the model has been trained. Further details for all participants, ROIs, other phases, and LLMs (OLMo-2, OLMo-0724, LLM-jp) can be found in Figures B.11B.16.
Figure 4: Probing results. a Evolution of probing accuracy for English/Japanese MMLU throughout the training process, assessed at layers 5, 15, 25, and 30 of OLMo-2. The horizontal axis indicates the probing accuracy. The vertical axis indicates the number of neurons that fall within each 0.01 accuracy bin. The legends corresponds to the number of training tokens. b Relationship between probing accuracy in the neurons of OLMo-2 across English MMLU, CSQA, and ARC. Each axis denotes the probing accuracy for the respective task, and the color gradient reflects the number of training tokens. The legend shows the correlation coefficient between certain two tasks.
@@
We now investigate how the neuron-level internal representations evolve over the course of training. For this purpose, we examine changes in the distribution of probing accuracy for each neuron at the early, middle, and final checkpoints (105B, 1259B, 3896B training tokens in OLMo-2) of Phase 3. Figure 4a presents the results for neurons in layers 5, 15, 25, and 30 of OLMo-2 when using MMLU. From these results, we find that, in the English MMLU setting, later-layer neurons progressively increase in probing accuracy, indicating that, as the LLMs acquire downstream task proficiency, they also develop more specialized neuron activations. Figures B.17B.19 show the results corresponding to all downstream tasks for OLMo-2, OLMo-0724, and LLM-jp, confirming that a similar trend emerges across every downstream task when each model is provided with the language on which it has been trained.
@@ 4.3 Changes in the Nature of Activations
As shown in Figure 5a, within-phase activations are highly similar, but substantial changes emerge at each phase transition. Interestingly, although the brain encoding accuracy does not vary dramatically between Phases 2 and 3, the underlying activations differ considerably. Next, we quantify the dimensionality of these activations using the IDs defined by GRIDE. Figure 5b shows that there is a strong similarity between the brain encoding accuracy and IDs during Phases 1 and 2, consistent with earlier findings that single-checkpoint encoding accuracy and dimensionality are highly correlated (Cheng & Antonello, 2024). Our findings extend this result by demonstrating a similarly robust correlation across multiple checkpoints. Nevertheless, the brain encoding accuracy and dimensionality are not perfectly congruent (see Phase 3), leaving open questions about the aspects of activation changes that intrinsic dimensionality alone does not capture.
@@ 5 Discussion and Conclusions
In this study, we interpreted the learning dynamics of LLMs from three perspectives: alignment between the brain and the LLMs, internal representations associated with downstream tasks, and actual downstream task performance. Our novel approach elucidated a three-stage phase transition—including the precise timing of each transition—that emerges consistently during training, even among LLMs that diverge substantially in both training data and architecture.
@@
Additionally, the learning dynamics that we have revealed are reminiscent of the “synaptic overproduction and pruning” posited in neuroscience, where an initial synaptic overgrowth eventually refines the neural circuits (Peter R., 1979; Bourgeois et al., 1994). These transient changes in the brain suggest the potential for more efficient neural information processing, leading to enchanced cognitive function. Thus, a promising avenue for future work involves examining whether the phenomenon observed in this study arises from the LLM pruning superfluous internal representations in pursuit of more refined ones.
@@
By capturing shifts in the activations of LLM neurons pertaining to downstream tasks, we elucidated how task-specific neurons form and how much specialization they share among tasks. These findings align with previous studies showing neurons that encode particular linguistic information (Tenney et al., 2019; Dai et al., 2022; Wang et al., 2022; Gurnee et al., 2023). We observed that neurons contributing substantially to certain tasks emerge abruptly in later layers as training progresses. Furthermore, the variations in neuron sharing across tasks suggest that differences in activation patterns—reflecting task-specific characteristics (e.g., required capabilities, difficulty levels, and answer formats)—are indeed discernible at the neuronal level. Taken together, our results offer a more comprehensive view of LLM neurons during the learning process, building on past work that focused solely on trained models. Further in-depth analyses of these neurons would be a fascinating avenue for future research.
@@ A.1 Large Language Models
We used the allenai/OLMo-2-1124-7B, allenai/OLMo-7B-0724-hf, llm-jp/llm-jp-3-7.2b, and LLM360/Amber models available on Hugging Face for OLMo-2, OLMo-0724, LLM-jp, and Amber. All checkpoints of LLM-jp will be made publicly available, although it has only released its final checkpoint. Tables 1 and 2 present an overview of the LLMs and detailed information on their respective training checkpoints used in this study. In selecting these checkpoints, we took particular care to ensure that the number of training tokens was as closely aligned as possible across the four LLMs.
Model Layers Width Params. Trn. Tokens (Ckpts.)
OLMo-2 32 4096 7.3B 1B - 3896B (28)
OLMo-0724 32 4096 6.89B 4B - 2724B (23)
LLM-jp 32 4096 7.29B 4.2M - 1258B (27)
Amber 32 4096 6.74B 3.5B - 1259B (18)
@@
MRI data were acquired using a 3T MAGNETOM Vida scanner (Siemens, Germany) with a standard Siemens 64-channel volume coil. Functional brain images based on the blood oxygenaction level-dependent (BOLD) signal were collected via a multiband gradient echo-planar imaging sequence (Moeller et al., 2010) (TR = 1,000 ms, TE = 30 ms, flip angle = 60, voxel size = 2 $\times$ 2 $\times$ 2 mm3, matrix size = 96 $\times$ 96, 72 slices with a thickness of 2 mm, slice gap 0 mm, FOV = 192 $\times$ 192 mm2, bandwidth 1736 Hz/pixel, partial Fourier 6/8, multiband acceleration factor 6). Anatomical data were acquired using the same 3T scanner using T1-weighted MPRAGE (TR = 2530 ms, TE = 3.26 ms, flip angle = 9, voxel size = 1 $\times$ 1 $\times$ 1 mm3, FOV = 256 $\times$ 256 mm2). The preprocessing of the fMRI data included motion correction, coregistration, and detrending. All participants provided written informed consent for this study, which was conducted under the approval of the relevant ethics and safety committee.

### v2

@@ 1 Introduction
Figure 1: Overview of the study. a Brain encoding analysis. b Probing analysis (top) and benchmark analysis (bottom). c Three phase-transition phenomena during the learning process of LLMs, as identified through the results of the encoding, probing, and benchmark analyses. Red, green, and blue lines indicate encoding, probing, and benchmark accuracies, respectively, across all LLM checkpoints.
@@ 3.1 Large language models
We analyze the learning dynamics of LLMs from three distinct perspectives. For this, we use four pre-trained models with available training checkpoints: OLMo-2 (OLMo et al., 2024), OLMo-0724 (Groeneveld et al., 2024), and LLM-jp (LLM-jp et al., 2024) for the main analysis, and Amber (Liu et al., 2023) for the control analysis. The number of checkpoints used for each model ranges from 18 to 28. Table 1 presents an overview of these LLMs. OLMo-2, OLMo-0724, and Amber are English-centric LLMs trained on publicly accessible datasets, while LLM-jp is a bilingual Japanese-English model trained on a roughly equal mix of Japanese and English. Each model uses a different training corpus and tokenizer, with vocabulary sizes ranging from 32,000 to 100,352. These LLMs have between 6.74–7.3B parameters, 32 hidden layers, and a hidden dimension of 4,096. They are all based on a decoder-only Transformer architecture (Vaswani et al., 2017), but each LLM incorporates a few critical modifications. There are also notable differences across these LLMs in terms of their layer normalization and attention mechanisms. See Section A.1 for further details on the models used. The multilayer perceptron (MLP) layers require most parameters and represent essential features (Bereska & Gavves, 2024; Geva et al., 2021). Thus, we use their neural activation as the internal representations of each model.
@@ 3.2 Brain encoding models
Our initial approach for analyzing the learning dynamics of the LLMs involves an investigation of how their activations progressively align with brain activity during the training process. Specifically, for each checkpoint described in Sections 3.1 and A.1, we perform an encoding analysis by evaluating the prediction accuracy of a learned linear mapping from each layer’s activations to brain activity (Naselaris et al., 2011; Nishimoto et al., 2011; Huth et al., 2012) (Figure 1a). These analyses are conducted separately for each participant.
@@
We use the Narrative Movie fMRI Dataset (Yamaguchi et al., 2024; Nakagi et al., 2024) for our brain encoding analysis. This dataset provides brain activity data from six healthy participants with normal or corrected-normal vision (three females; ages 22–40, mean = 28.7), while they freely watched 8.3 hours of movies or drama series inside a 3T functional Magnetic Resonance Imaging (fMRI) scanner. All participants were native Japanese speakers. The dataset includes nine video clips of movies or drama series as stimuli: eight international titles and one Japanese title. The international titles were dubbed into Japanese, allowing the participants to view all clips in Japanese. The dataset also contains three types of natural language annotation from the videos. We use the Narrative Content annotation, which describes the background story of the scene at 5-s intervals, for the main analysis and the Objective Information annotation, which describes the objects in the scene every second, for the control analysis, both in English and Japanese. See Section A.2 for additional dataset details. In this study, we use 29,993 seconds of data covering all six participants. We divide the data into training and test datasets. All prediction performance results are computed using the test dataset. Specifically, we use the fMRI scanning sessions corresponding to the last split of each movie or drama series (7,737 seconds in total) as the test data. The remainder of the sessions (22,262 seconds in total) forms the training data.
@@
We extract the activations of the LLMs for the annotations from each hidden layer. They consist of several tokens for each time point. We then average the activations across tokens. Because multiple annotations exist for each second, we extract the activations for each annotator and then average them across all annotators. We then train an L2-regularized linear model to predict the voxel-level brain activity from the corresponding activations of the LLM neurons. We estimate the model weights from the training data, then apply them to the test data. The regularization parameters are determined for each voxel by cross-validation during training. We then evaluate the prediction accuracy by computing the Pearson’s correlation coefficients between the predicted and measured fMRI signals. Statistical significance is assessed using a blockwise permutation test that compares the correlations between predicted and measured signals against the correlations obtained after shuffling the measured signals. We set the threshold for statistical significance to $p<0.05$ and correct for multiple comparisons using the FDR procedure. We model the hemodynamic delay in the BOLD signal, assumed to be 8–10 seconds. See Section A.3 for additional details about model construction and region-of-interest (ROI) analyses. In the main analysis, we focus on later layers within each LLM. Specifically, we focus on layer 25 in both OLMo-2 and LLM-jp, and layer 30 in OLMo-0724, which show the greatest checkpoint-wise changes and are strongly involved in each model’s emergent behavior across all evaluations (encoding, probing, and benchmark analyses; see Section A.6 for details). We confirm that the layers surrounding the main focal layer exhibit similar patterns of results (see Figure B.5).
@@ 3.3 Probing
As the second approach, we investigate how downstream task-relevant representations are progressively acquired within the LLMs during the training process. To this end, we perform a probing analysis by evaluating how well linear models can predict neuron activations from the answer labels of downstream tasks across all hidden layers (Figure 1b, top). These analyses are conducted separately for each LLM checkpoint.
@@
We use four downstream tasks: Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2020), CommonsenseQA (CSQA) (Talmor et al., 2019), AI2 Reasoning Challenge (ARC) (Clark et al., 2018), and HellaSwag (Zellers et al., 2019). For both the probing and benchmark analyses (see Section 3.4), we use 5-shot prompts in both English and Japanese. We exclude prompts exceeding each model’s maximum context length. See Section A.4 for additional details about downstream datasets and Section A.7 for examples of the prompts used in these datasets. In the probing analysis, each dataset is split into training and test datasets at a 4:1 ratio. All prediction performance results are computed using the test dataset.
@@
We first feed the downstream task prompts into the LLMs and extract the final-token neuron activations in each layer. For each sample in the downstream tasks, we create an answer matrix (the number of samples $\times$ the number of choices), where the correct and incorrect choices are labeled as 1 and 0, respectively. Our objective is to learn a mapping from this answer matrix to the activations across all layers of the LLMs. To this end, we use L2-regularized linear regression to estimate the linear weights that transform the answer matrix corresponding to the training data into the observed activations in the LLMs, and then apply these learned weights to the test data. The regularization parameter is optimized via 4-fold cross-validation for each neuron in the training dataset. We then evaluate the prediction accuracy by computing the Pearson’s correlation coefficients between the predicted and actual activations. The amount of information necessary for each task that is retained at each LLM checkpoint is quantitatively evaluated at the neuron level. In this study, we focus on the same layers as in the encoding analysis (Section 3.2).
@@ 4.1 Learning dynamics of LLMs
Figure 3: Changes in the relationship with the brain. Differences in encoding accuracy among checkpoints for three participants (DM06, DM03, and DM07), projected onto the inflated (top, lateral, and medial views) and flattened cortical surface (occipital areas are at the center, only for DM06), for both the left and right hemispheres. We have chosen OLMo-2’s checkpoints that capture the transitions at each phase of the learning dynamics. Brain regions with significant accuracy at either of the two checkpoints are colored ($p<0.05$, FDR-corrected). Voxels exhibiting higher accuracy at the later checkpoint are indicated in red, whereas those exhibiting higher accuracy at the earlier checkpoint are indicated in blue.
@@ 4.2 What happens during each phase?
Overall, these findings indicate that once the LLMs begin following the task instructions (Phase 1), their internal representations exhibit better global alignment with brain activity. During the subsequent stagnation phase (Phase 2), this global alignment partially recedes, and after Phase 3, the LLMs realign with the brain, reinforcing the similarity to brain regions involved in semantic processing, in particular. Notably, these patterns only emerge when the input language matches one on which the model has been trained. Further details for all participants, ROIs, other phases, and LLMs can be found in Figures B.11B.16.
Figure 4: Probing results. a Evolution of probing accuracy for English/Japanese MMLU throughout the training process, assessed at layers 5, 15, 25, and 30 of OLMo-2. The horizontal axis indicates the probing accuracy. The vertical axis indicates the number of neurons that fall within each 0.01 accuracy bin. The legends corresponds to the number of training tokens. b Relationship between probing accuracy in the neurons of OLMo-2 (layer 25) across English MMLU, CSQA, and ARC. Each axis denotes the probing accuracy for the respective task, and the color gradient reflects the number of training tokens. The legend shows the correlation coefficient between certain two tasks.
@@
We now investigate how the neuron-level internal representations evolve over the course of training. For this purpose, we examine changes in the distribution of probing accuracy for each neuron at the early, middle, and final checkpoints (302B, 1259B, 3896B training tokens in OLMo-2) of Phase 3. Figure 4a presents the results for neurons in layers 5, 15, 25, and 30 of OLMo-2 when using MMLU. From these results, we find that, in the English MMLU setting, later-layer neurons progressively increase in probing accuracy, indicating that, as the LLMs acquire downstream task proficiency, they also develop more specialized neuron activations. Figures B.17B.19 show the results corresponding to all downstream tasks for OLMo-2, OLMo-0724, and LLM-jp, confirming that a similar trend emerges across every downstream task when each model is provided with the language on which it has been trained.
@@ 4.3 Changes in the nature of activations
As shown in Figure 5a, within-phase activations are highly similar, but substantial changes emerge at each phase transition. Interestingly, although the brain encoding accuracy does not vary dramatically between Phases 2 and 3, the underlying activations differ considerably. Next, we quantify the dimensionality of these activations using the IDs defined by GRIDE. Figure 5b shows that there is a strong similarity between the brain encoding accuracy and IDs during Phases 1 and 2, consistent with earlier findings that single-checkpoint encoding accuracy and IDs are highly correlated (Cheng & Antonello, 2024). Our findings extend this result by demonstrating a similarly robust correlation across multiple checkpoints. Nevertheless, the brain encoding accuracy and IDs are not perfectly congruent (see Phase 3), leaving open questions about the aspects of activation changes that intrinsic dimensionality alone does not capture.
@@ 5 Discussion and conclusions
In this study, we interpreted the learning dynamics of LLMs from three perspectives: alignment between LLMs and the brain, internal states associated with downstream tasks, and downstream task performance. Our novel approach elucidated a three-stage phase transition—including the precise timing of each transition—that emerges consistently during training, even among LLMs that diverge substantially in both training data and architecture.
@@
Additionally, the learning dynamics that we have revealed are reminiscent of the “synaptic overproduction and pruning” posited in neuroscience, where an initial synaptic overgrowth eventually refines the neural circuits (Peter R., 1979; Bourgeois et al., 1994). These transient changes in the brain suggest the potential for more efficient neural information processing, leading to enhanced cognitive function. Thus, a promising avenue for future work involves examining whether the phenomenon observed in this study arises from the LLM pruning superfluous internal representations in pursuit of more refined ones.
@@
By capturing shifts in LLM neuron activations related to downstream tasks, we elucidated how task-specific neurons form and how much specialization they share among tasks. These findings align with previous studies showing neurons that encode particular linguistic information (Tenney et al., 2019; Dai et al., 2022; Wang et al., 2022; Gurnee et al., 2023). We observed that neurons contributing substantially to certain tasks emerge abruptly in later layers as training progresses. Moreover, the variations in neuron sharing across tasks suggest that differences in activation patterns—reflecting task-specific characteristics (e.g., required capabilities, difficulty levels, and answer formats)—are indeed discernible at the neuronal level. Taken together, our results offer a more comprehensive view of LLM neurons during the learning process, building on past work that focused solely on trained models. Further in-depth analyses of these neurons would be a fascinating avenue for future research.
@@ A.1 Large language models
We used the allenai/OLMo-2-1124-7B, allenai/OLMo-7B-0724-hf, llm-jp/llm-jp-3-7.2b, and LLM360/Amber models available on Hugging Face for OLMo-2, OLMo-0724, LLM-jp, and Amber. All checkpoints of LLM-jp will be made publicly available, although it has only released its final checkpoint. Tables 1 and 2 present an overview of the LLMs and detailed information on their respective training checkpoints used in this study. We used 28 checkpoints for OLMo-2 (1B-3,896B training tokens), 23 for OLMo-0724 (4B–2,724B), 27 for LLM-jp (4.2M–1,258B), and 18 for Amber (3.5B–1,259B). In selecting these checkpoints, we took particular care to ensure that the number of training tokens was as closely aligned as possible across the four LLMs.
Model Layers Width Params. Vocab. sizes Trn. Tokens (Ckpts.)
OLMo-2 32 4096 7.3B 100352 1B - 3896B (28)
OLMo-0724 32 4096 6.89B 50304 4B - 2724B (23)
LLM-jp 32 4096 7.29B 99584 4.2M - 1258B (27)
Amber 32 4096 6.74B 32000 3.5B - 1259B (18)
@@
MRI data were acquired using a 3T MAGNETOM Vida scanner (Siemens, Germany) with a standard Siemens 64-channel volume coil. Functional brain images based on the blood oxygenation level-dependent (BOLD) signal were collected via a multiband gradient echo-planar imaging sequence (Moeller et al., 2010) (TR = 1,000 ms, TE = 30 ms, flip angle = 60, voxel size = 2 $\times$ 2 $\times$ 2 mm3, matrix size = 96 $\times$ 96, 72 slices with a thickness of 2 mm, slice gap 0 mm, FOV = 192 $\times$ 192 mm2, bandwidth 1736 Hz/pixel, partial Fourier 6/8, multiband acceleration factor 6). Anatomical data were acquired using the same 3T scanner using T1-weighted MPRAGE (TR = 2530 ms, TE = 3.26 ms, flip angle = 9, voxel size = 1 $\times$ 1 $\times$ 1 mm3, FOV = 256 $\times$ 256 mm2). The preprocessing of the fMRI data included motion correction, coregistration, and detrending. All participants are right-handed, native Japanese speakers and provided written informed consent for this study, which was conducted under the approval of the relevant ethics and safety committee.